LABORATORY INVESTIGATION OF THE AIRGLOW BANDS

\underline{BRIAN DROUIN}, SHANSHAN YU, TIMOTHY J. CRAWFORD, CHARLES E. MILLER, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109; JENG-HWA YEE, Applied Physics Laboratory, The Johns Hopkins University, 111000, Johns Hopkins Rd, Laurel, MD 20723-6099.

We report the first high spectral resolution laboratory measurements of oxygen A-band night glow simulated using a static discharge cell. Our static discharge system reproduces the conditions of the mesospheric oxygen night glow - suggesting O(^{1}D) + O_{2} as the primary source of the emission. Additionally, use of the static cell has enabled us to collect spectra for rare molecular oxygen isotopologues using isotopically enriched samples. The (0,0), (0,1), and (1,1) b − Xvibrational bands were observed with a Bruker 125 HR for all six isotopologues. The (1,2) and (2,2) bands were observed also for the main isotopologue. The frequencies of the observed (0,1) transitions resolved discrepancies in Raman data for (16-17, 17-17, and 17-18), enabling us to improve the vibrational parameterization of the ground electronic state in the global fit of Yu et al.^{a} Rotationally resolved intensities were determined for the (0,0), (0,1) and (1,1) bands. The experimental band intensity ratios I(0,0)/I(0,1) = 13.6 and I(0,0)/I(1,1) = 60 are in excellent agreement with the recent mesospheric remote sensing data.

aS. Yu, C.E. Miller, B.J. Drouin, H.S.P. Müller, J. Chem. Phys. 136, 024304, 2012